# Black-hole spectroscopy from a giant quantum vortex

### v1

@@
In panels b and d, we evaluate Eq. (5) and plot $-\log_{10}(\mathrm{Res})$ in the complex frequency plane, aligning the real-axis extent with the orange-shaded regions in panels a and c. Solutions to the resonance condition, $\mathrm{Res}(\omega)=0$, appear as local maxima, marking the corresponding QNM frequencies. Modes below the light-ring frequency (white dashed line) are also known as quasibound states (white circles) and correspond to minimally-damped waves trapped between the outer boundary and the finite potential barrier. In contrast, the damping of QNMs above the light-ring frequency (white crosses) is more significant. To illustrate how these resonances relate to the QNMs of an open system, we track their migration as the outer boundary’s reflectivity smoothly varies from 0 to 1 [10]. Notably, all QNMs shift toward the real frequency axis, indicating a reduction in their decay rates. Quasibound states can be particularly identified as higher overtones of the open system that rapidly shift to lower real and nearly vanishing imaginary frequencies.
The obtained QNM frequencies are compared with the interface noise spectra in panels a and c, where they are marked by vertical green lines. The agreement between the observed and predicted excitations confirms the validity of our model and the chosen boundary condition. QNMs appear in the spectra as distinct peaks. For $m=-7$ (panel a), peaks corresponding to the quasibound state (located at $6.9$ Hz) and the fundamental QNM ($8.0$ Hz), are well-defined because their low damping rate corresponds to narrow peak widths. In contrast, the higher-frequency overtone ($8.5$ Hz) appears as a broader peak due to its faster decay. For $m=-12$ (panel c), the experimental spectrum reveals an additional mode within the region of interest. Here, the first two quasibound states again exhibit low damping, producing narrow peaks, while other QNMs appear smeared out due to their higher dissipation rates.
@@
Excitations above the orange-shaded regions display an additional set of peaks. These high-momentum waves only minimally interact with the effective potential and propagate between the giant vortex core and the boundary. Their dynamics are primarily influenced by interactions with the giant vortex itself. As vorticity dominates the flow inside the vortex core, accurately modelling these waves warrants further enquiries.
@@
Panel e provides a more comprehensive view of the rich noise spectrum. It presents a full wave spectrogram, including co-rotating waves ($m>0$ bands). The structure in this half-plane of the Fourier space reveals bound states, which are standing waves confined between the potential barrier (for co-rotating waves, the potential diverges at small radii) and the boundary [9]. In the opposite half-plane, a broad band of excitations (brighter tones, excluding the $m=-1$ band, which is an experimental artifact) emerges across a wide range of frequencies and azimuthal numbers. A closer examination is provided in panel f, where each azimuthal band is normalised by its mean amplitude, highlighting strong excitations matching the predicted quasibound state frequencies (white circles) and other QNM frequencies (white crosses), particularly for $m\gtrsim-10$. For $m\lesssim-10$, low-frequency modes are overshadowed by broad, higher-frequency excitations. These waves corroborate theoretical estimates of additional resonant modes (white circles and crosses), both above and below the light-ring frequency (solid white line). As higher-frequency QNMs exhibit the largest damping, their Fourier peaks acquire a finite width, limiting the resolution of individual modes in this part of the spectrum.
@@
Recent studies [41, 42] suggest that detecting spectral modifications in binary black hole mergers can provide valuable insights into their environments. Such observations enable detailed studies of spectral stability in black holes surrounded by plasma [43], ultralight bosonic fields [44, 45], or those in asymptotically anti-de Sitter spacetimes [46, 47, 48] where various boundary conditions have been explored. However, direct detection remains challenging, as the environmental effects may be too subtle for current or next-generation gravitational-wave detectors to resolve [41].
@@
By confirming the predicted occurrence of multiple quasinormal modes (QNMs), our experimental platform provides compelling proof of principle, demonstrating the critical role of spatial confinement in shaping the excitation spectrum. As a next step, gravity simulators based on wave dynamics in quantum liquids will provide new opportunities to harness noise spectroscopy in bounded systems. Techniques developed for gravitational-wave analysis can be adapted to laboratory investigations, revealing additional overtones and serve as a test bed for data analysis techniques. Moreover, by refining the energy injection mechanism, whether through narrowband mechanical vibrations or controlled variations in the flow drive, we could selectively amplify specific QNMs. This process would push the system beyond the linear perturbation regime, allowing for the investigation of primary resonances scattering into an expanding family of secondary resonances [49], offering valuable insights into QNM interactions [50]. The synergy between laboratory analogues and astrophysical black holes fosters a virtuous cycle, where insights from controlled experiments feed back into our modelling of astrophysical black holes, and vice versa.
@@ Acknowledgements
We are grateful to Sebastian H. Völkel for valuable feedback and fruitful discussions. P.S., P.Š., S.P., C.F.B., R.G. and S.W. extend their appreciation to the Science and Technology Facilities Council for their generous support within Quantum Simulators for Fundamental Physics (ST/T006900/1, ST/T005858/1, and ST/T00584X/1), as part of the UKRI Quantum Technologies for Fundamental Physics program. L.S. and S.W. gratefully acknowledge the support of the Leverhulme Research Leadership Award (RL-2019-020). M.R. acknowledges partial support from the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, Brazil, grant 315991/2023-2), and from the São Paulo Research Foundation (FAPESP, Brazil, grant 2022/08335-0). S.W. also acknowledges the Royal Society University Research Fellowship (UF120112). R.G. and S.W. acknowledge support from the Perimeter Institute. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Research, Innovation and Science.

### v2

@@
In panels b and d, we evaluate Eq. (5) and plot $-\log_{10}(\mathrm{Res})$ in the complex frequency plane, aligning the real-axis extent with the orange-shaded regions in panels a and c. Solutions to the resonance condition, $\mathrm{Res}(\omega)=0$, appear as local maxima, marking the corresponding QNM frequencies. The imaginary part of these maxima is related to the intrinsic damping of the QNMs due to the presence of the absorbing boundary at the vortex core (see Methods for a comparison with the viscous damping rate). Quasibound states (white circles), i.e. modes below the light-ring frequency (white dashed line), correspond to waves trapped between the outer boundary and the finite potential barrier. As expected, these waves are only minimally damped. In contrast, the damping of QNMs above the light-ring frequency (white crosses) is more significant, as they are capable of propagating over the barrier and eventually leaving the system through the vortex core. To illustrate how these resonances relate to the QNMs of an open system, we compute their trajectories in the complex frequency plane as the reflectivity of the outer boundary smoothly increases from 0 to 1 [10]. Notably, all QNMs shift towards the real frequency axis, indicating a reduction in their decay rates. Quasibound states can be particularly identified as higher overtones of the open system that rapidly shift to lower real and nearly vanishing imaginary frequencies.
The obtained QNM frequencies are compared with the experimental spectra in panels a and c. These predictions are indicated by green vertical bands, whose thickness corresponds to statistical error due to uncertainties in the flow parameters $C$ and $\Omega$. Close agreement of our predictions and distinct excitations in the experimental spectra confirms the validity of the theoretical model and validate the chosen boundary condition. For $m=-4$ (panel a), the quasibound state, located at $5.6$ Hz, appears as a well-defined peak due to its low damping rate. Another resonance, now at $6.9$ Hz, lies close to a multiple of the driving frequency, but the notably large amplitude of its spectral peak, greater than those of previous drive harmonics, strongly suggests that the predicted resonance is indeed realised in the experiment. For $m=-16$ (panel c), the first quasibound state is less prominent, as it is overshadowed by higher-frequency resonances. These resonances appear as broader, overlapping peaks, associated with stronger damping, as is indicated in panel d by their position along the imaginary axis.
@@
Note that excitations above the orange-shaded regions display additional peaks. These high-momentum waves only minimally interact with the effective potential and propagate between the giant vortex core and the boundary. Their dynamics are primarily influenced by interactions with the giant vortex itself and can potentially include quadratic QNMs [40]. As vorticity dominates the flow inside the vortex core, accurately modelling these waves warrants further enquiries.
@@
Panel e provides a more comprehensive view of the rich noise spectrum. It presents a full wave spectrogram, including co-rotating waves ($m>0$ bands). The structure in this half-plane of Fourier space reveals bound states, which are standing waves confined between the boundary and potential barrier that diverges at small radii [9]. In the opposite half-plane, a broad band of excitations (brighter tones, excluding the $m=-1$ band, which is an experimental artifact) emerges across a wide range of frequencies and azimuthal numbers. A closer examination is provided in panel f, where each azimuthal band is normalised by its mean amplitude, highlighting strong excitations matching the predicted quasibound state frequencies (white circles) and other QNM frequencies (white crosses), particularly for $m\gtrsim-8$. For $m\lesssim-8$, low-frequency modes are overshadowed by broad, higher-frequency excitations. These waves corroborate theoretical estimates both above and below the light-ring frequency (dashed white line). As higher-frequency QNMs exhibit the largest damping, their Fourier peaks acquire a finite width, limiting the resolution of individual modes in this part of the spectrum.
@@
Recent studies [41, 42] suggest that spectral modifications in the ringdown emitted after a binary black hole merger provide information about the environment of the resulting black hole. Such insights motivate spectral stability studies in black holes surrounded by plasma [43], ultralight bosonic fields [44, 45], or those in asymptotically anti-de Sitter spacetimes [46, 47, 48], where various boundary conditions are under scrutiny. However, spectra modified by realistic environments may be too subtle to be resolved by the current or next-generation gravitational-wave detectors [41].
@@
Within the framework of the black-hole spectroscopy programme, we confirm the predicted occurrence of multiple QNMs and infer the outer boundary condition from experimental data. As a next step, gravity simulators based on wave dynamics in quantum liquids will provide new opportunities to exploit noise-driven spectra in spatially bounded systems. Emerging techniques, developed for gravitational-wave analysis, can be benchmarked using laboratory experiments. Moreover, by refining the energy-injection mechanism, whether through narrowband mechanical vibrations or controlled modulation of the flow drive, we can selectively amplify specific QNMs. This approach pushes the system beyond the linear perturbation regime and enables the exploration of primary resonances scattering into an expanding family of secondary excitations [49]. By analysing QNM interactions [50], we seek to identify the conditions under which nonlinear effects dominate the QNM spectrum. The synergy between laboratory analogues and astrophysical black holes fosters a virtuous cycle, where insights from controlled experiments feed back into our modelling of astrophysical black holes, and vice versa.
@@ Acknowledgements
We are grateful to Cameron R. D. Bunney, Gregorio Carullo, Christoph Eigen, Grégoire Ithier, Maxime Jacquet, Friedrich König, Xavier Rojas, and Sebastian H. Völkel for valuable feedback and fruitful discussions. P.S., P.Š., S.P., C.F.B., R.G. and S.W. extend their appreciation to the Science and Technology Facilities Council for their generous support within Quantum Simulators for Fundamental Physics (ST/T006900/1, ST/T005858/1, and ST/T00584X/1), as part of the UKRI Quantum Technologies for Fundamental Physics programme. L.S. and S.W. gratefully acknowledge the support of the Leverhulme Research Leadership Award (RL-2019-020). M.R. acknowledges partial support from the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, Brazil, grant 315991/2023-2), and from the São Paulo Research Foundation (FAPESP, Brazil, grants 2022/08335-0 and 2025/02701-3). S.W. also acknowledges the Royal Society University Research Fellowship (UF120112). R.G. and S.W. acknowledge support from the Perimeter Institute. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Research, Innovation and Science.
Author Contributions
All authors contributed substantially to the work. Acquisition and analysis of experimental data by P.S., L.S., and P.Š. Theoretical modelling developed by L.S., S.P., and M.R. All authors contributed to the discussion of the results and the writing of the manuscript. Funding acquisition by S.W., R.G., and C.F.B. Overall conceptualization and project supervision by S.W.